A new car is purchased for 22300 dollars. The value of the car depreciates at 10.25% per year. What will the value of the car be, to the nearest cent, after 6 years?



1
Expert's answer
2022-03-08T10:02:03-0500

For annual compound interest

"R=P(1-0.1025)^6=22300*0.8975^6\approx11655" dollars value of the car will be after 6 years